OpenStudy (anonymous):

Write the equation of a line in slope intercept form that is parallel to the line y=1/3x=5 and passes through (-9, 5). Do not use any spaces. Slope should be a reduced improper fraction.

OpenStudy (anonymous):

y-y1=m(x-x1) m=slope (x1,y1)= point We know that parallel lines have the same slope so just find the slope of y=1/3x=5 And you are already given a point : (-9, 5) Plug them in and you ae done

OpenStudy (anonymous):

y=-9x+5 ?